Given a standard normal distribution with mean =105 and standard deviation = 28, find P(72 < X < 111)


1
Expert's answer
2022-02-02T14:04:19-0500

 P(72 < X < 111)

we use z calculations

z (72) = ( 72 - 105) / (28)

= -1.178571429

z (111) = ( 111 - 105 ) / (28)

= 0.214285714

using two decimal places ,we may define

P ( -1.18 < z < 0.21 )

using normal tables, we obtain

= ( 0.58317 - 0.11900 )

= 0.46417 which is the required solution.